Find the first and second derivative of the following:



1. y=cos x²


2.y=sinx cos x²


3.y=xsin x²


4.y=x cos x²

1
Expert's answer
2022-03-04T06:48:53-0500

"1) y=\cos x^2\\
y' =-\sin x^2\cdot 2x=-2x\sin x^2\\
y''=-2\sin x^2 - 2x\cdot\cos x^2\cdot2x=\\
=-2\sin x^2-4x^2 \cos x^2;\\
2) y=\sin x\cos x^2\\
y'=\cos x \cos x^2 - 2x\sin x \sin x^2\\
y''=-\sin x \cos x^2 -\cos x \sin x^2 \cdot 2x -\\
-\cos x \sin x^2 \cdot 2x - \sin x\cos x^2\cdot2x\cdot2x-\\
-\sin x \sin x^2 \cdot2=\\
=-\sin x\cos x^2 -4x \cos x \sin x^2-\\
-4x^2 \sin x\cos x^2 -2\sin x \sin x^2;\\
3) y=x \sin x^2\\
y'=\sin x^2 +x\cdot\cos x^2\cdot2x=\\
=\sin x^2 +2x^2 \cos x^2\\
y''=\cos x^2 \cdot 2x+4x \cos x^2 -2x^2 \sin x^2 \cdot2x=\\
=6x \cos x^2-4x^3 \sin x^2;\\
4) y=x \cos x^2\\
y'= \cos x^2 -x \sin x^2 \cdot 2x=\\
=\cos x^2 - 2x^2 \sin x^2\\
y''= -\sin x^2\cdot 2x- 4x \sin x^2-2x^2 \cos x^2\cdot2x=\\
=-6x \sin x^2-4x^3\cos x^2."